Word Problem
Two angles of a triangle measure 21 degrees and 86 degrees. What is the measure of the third angle?

Answer by wuwei96815(245)

The angles of a triangle always add up to 180 degrees.

So the answer is 180 - 21 - 86 = 180 - 107 = 73 degrees.